OpenStudy (yrivers36):

How do I do square root y^3 * cubed square root y^2? I am totally confused with this problem.

OpenStudy (anonymous):

is this \[y^{3}\sqrt[3]{y^2}\] you are working with?

OpenStudy (anonymous):

under a square root sign?

OpenStudy (anonymous):

If I'm reading you right, then your equation would look like y^(3/2)*y^(2/3). With same bases, you would add your exponents, so 3/2+2/3=13/6. Your answer is y^13/6, which you can simplify.

OpenStudy (anonymous):

Remember that roots are just the base raised to 1 over whatever you are rooting to. For example, the square root of y is the same as y^(1/2), and the cube root is y^(1/3). Prove this to yourself on a calculator with y as any number. If your base already has an exponent, that exponent becomes the numerator. So square root of y^3 is y^(3/2) and cube root of y^2 is y^(2/3).

OpenStudy (yrivers36):

Mfang its not an equation. Its radicals.

OpenStudy (yrivers36):

Daomowon its \[\sqrt{y ^{3}}\]

OpenStudy (yrivers36):

thats not what i typed...omg going crazy over this problem

OpenStudy (yrivers36):

its that prob X sqrt y^3

OpenStudy (yrivers36):

there is a sqrt sign with the y^3

OpenStudy (yrivers36):

mfang Im sorry your correct

OpenStudy (anonymous):

no need to worry, whatever I just typed didn't come up but \[y^3 = y^2*y\], then